15)Find the constants a and b such that the function f (z)=ax-3y+i(bx-2y) is analytic?

Explain the problem with step by step process?


1
Expert's answer
2021-09-24T02:12:59-0400

Given:

f (z)=ax-3y+i(bx-2y) 

We solve this as follows:

2x + y = 32 x + 3y = 36 21122 + 8 ≟ 32 12 + 3182 ≟ 36 32

Values of a and b are: 32 and 36

f (z)=ax-3y+i(bx-2y) is analytic